rxcklesskaisher:

5c+cd5, c, plus, c, d when c=\dfrac15c= 5 1 ​ c, equals, start fraction, 1, divided by, 5, end fraction and d=15d=15d, equals, 15.

AZ:

Can you take screenshots instead so I don't have to sit here trying to figure out what your question says? Like just look at what you're copy-pasting but I think that's saying 5c + cd and c = 1/5 and d= 15 and we're supposed to find the value of the expression when c = 1/5 and d = 15

AZ:

Am I correct in my assumption?

rxcklesskaisher:

AZ:

Ah much better so all you have to do is plug in \(\color{red}{c} = \color{red}{\dfrac{1}{5}}\) and \(\color{orange}{d} = \color{orange}{15}\) so put the numbers where the letters are and solve \(5\color{red}{c} + \color{red}{c}\color{orange}{d}\) \(5\left(\color{red}{\dfrac{1}{5}} \right)+ \left(\color{red}{\dfrac{1}{5}} \right)\color{orange}{15}\) so first what is 5 * 1/5 and then what is 1/5 * 15

rxcklesskaisher:

18

AZ:

how did you get that first, what is 5 * 1/5 =

rxcklesskaisher:

o.2

rxcklesskaisher:

@AZ

AZ:

can you use a calculator? 5 * 1/5 is the same thing as 5 / 5 what is 5 divided by 5 =

rxcklesskaisher:

1

AZ:

okay good and now can you do 1/5 * 15 it's the same thing as 15 * 1/5 which is 15 / 5 =

rxcklesskaisher:

ok

AZ:

what do you get what is 15 / 5 =

rxcklesskaisher:

3

AZ:

good so now we went from \(\color{red}{c} = \color{red}{\dfrac{1}{5}}\) and \(\color{orange}{d} = \color{orange}{15}\) \(5\color{red}{c} + \color{red}{c}\color{orange}{d}\) \(5\left(\color{red}{\dfrac{1}{5}} \right)+ \left(\color{red}{\dfrac{1}{5}} \right)\color{orange}{15}\) \(1+ \left(\color{red}{\dfrac{1}{5}} \right)\color{orange}{15}\) \(1+ 3\) so what is 1 + 3

rxcklesskaisher:

3

AZ:

one plus three is what? 1 + 3 = ?

rxcklesskaisher:

its looked like division sorry lolllll

rxcklesskaisher:

4

AZ:

and that's your final answer

rxcklesskaisher:

oh thanks again

rxcklesskaisher:

ur a pro

AZ:

No problem!

AZ:

Hopefully you were able to understand what we did :)